Suppose the GDP per capita (income) has increased from $5000 to $10,000, whereas the demand for luxury apartments has increased from 1000 to 2000, whereas the demand for burgers has increased from 10,000 to 12,000. Calculate the income elasticity of demand for apartments and burgers



1
Expert's answer
2021-11-28T17:56:05-0500

Solution:

Income elasticity for demand (YED) = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; income}"

 

Income elasticity for demand (YED) = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{I_{2} -I_{1}}{(I_{2}+I_{1})/2 }"

 

Income elasticity for demand (YED) for apartments:

Q1 = 1000                                  I1 = 5,000

Q2 = 2000                                  I2 = 10,000


Income elasticity for demand (YED) = "\frac{2000 -1000}{(2000+1000)/2 } \div \frac{10,000 -5,000}{(10,000+5,000)/2 } = \frac{1000}{1500} \div\frac{5,000}{7,500} = \frac{0.67}{0.67} = 1"


Income elasticity for demand (YED) for apartments = 1


Income elasticity for demand (YED) for burgers:

Q1 = 10000                                  I1 = 5,000

Q2 = 12000                                  I2 = 10,000


Income elasticity for demand (YED) = "\frac{12000 -10000}{(12000+10000)/2 } \div \frac{10,000 -5,000}{(10,000+5,000)/2 } = \frac{2000}{11000} \div\frac{5,000}{7,500} = \frac{0.18}{0.67} = 0.27"


Income elasticity for demand (YED) for burgers = 0.27